OpenStudy (anonymous):

Verify the identities (1+sinx)(1-sinx)=cos^2x

OpenStudy (psymon):

\[(a+b)(a-b) = a ^{2}-b ^{2}\] Multiply it out and then use the pythagorean identity: \[\sin ^{2}x + \cos ^{2}x = 1\]

OpenStudy (mathstudent55):

Multiply out the left side. It's the product of a sum and a difference. What do you get?